Two angles are complementary. One contains 30° more than the other. Find both angles.
Answer by pie90010(44)

let angle 1 be x
let angle 2 be y
since they are complimentary, they add up to 90, so we know
x + y = 90°
since one contains 30° more, we know
x = y + 30°
line them up
x - y = 30°
x + y = 90
the y's cancel
2x = 120
x = 60
now that you have x, plug it into the original equation to get y
x + y = 90
y = 30
the answer is the 2 angles are
30, and 60